# SPIN-Bench: How Well Do LLMs Plan Strategically and Reason Socially?

### v1

@@ 4.1 Outline and Setup
We evaluate a suite of popular LLMs that encompasses both closed-source commercial and open-source systems. The closed-source models include GPT-4o (Hurst et al., 2024), GPT-4o mini (OpenAI, 2024a), o1 (Jaech et al., 2024), o1-preview (OpenAI, 2024c), o1-mini (OpenAI, 2024b), o3-mini (OpenAI, 2025), GPT-4-turbo, GPT-3.5-turbo, Claude 3.5 Sonnet (Anthropic, 2024a), and Claude 3.5 Haiku (Anthropic, 2024b). Our open-source models include DeepSeek-R1 (Guo et al., 2025), Llama3-70b, Llama3.1-70b, Llama3.2-3b, Llama3.3-70b (AI, 2024a, b, c, d), Qwen2.5-72b (Team, 2024), and Mistral-7b (Jiang et al., 2023). Depending on the task, we select representative subsets to manage computational costs; for instance, classical planning experiments focus on nine well-known models (o1, o1-mini, o3-mini, GPT-4o, GPT-4-turbo, Claude 3.5 Sonnet, Claude 3.5 Haiku, Deepseek R1 and Llama3.3-70b).
@@
We adopt multiple metrics suitable for each domain. In classical planning tasks, we use accuracy, N-Step Look Ahead, and plan validation through VAL (Planning, 2019). Competitive board games are tested via LLM vs. solver and LLM vs. LLM matches, Top-3 action percentages and we maintain internal Elo scores for leaderboards (Appendix B.1). In cooperative and strategic settings, we measure final scores, success rates for different action types, and advanced LLM-assisted negotiation metrics when negotiations are enabled. Details of experimental settings, including prompt templates and domain rules, appear in Appendix A.
@@ 4.2 Overall Model Performance Across Tasks
As summarized in Tables 1, 2, and further detailed in Appendix B.1, the strongest closed-source models (e.g., o1, o1-mini, o3-mini) generally lead in classical planning, board-game play, and cooperative scenarios. Among open-source models, DeepSeek-R1 and higher-parameter Llama3 variants tend to outperform smaller-scale systems, although their margins behind the top-tier commercial LLMs remain significant.
@@ 4.3.1 Spatial Reasoning
To investigate whether LLMs’ planning deficits stem from weaker spatial understanding, we designed tasks requiring each model to track positions across sequences of relative movements. In these tasks, the LLM is given an initial spatial coordinate (e.g., (2,2)) along with a series of movements(up, down, left, right), and it tb must produce the final coordinate. We evaluated all models on 90 scenes, each with a varying number of steps, to assess their spatial reasoning capabilities. Figure 2 plots the accuracy of each model against the length of the movement trajectory. Notably, o1-mini and GPT-4o exhibit declining performance as the number of steps increases, whereas o1 sustains perfect accuracy (100%) up to 29 steps.
Spatial Orders in Diplomacy.
We further examine spatial understanding in a more strategic setting, Diplomacy, where orders like Move and Attack require consistent tracking of boards and adjacency relations. Table 2 shows that o1 surpasses other models by on success rate of basic orders. However, all models’ performance drops markedly for complex moves requiring multi-step or multi-agent logic (e.g., Support-Self and Support-Others), suggesting that deeper spatial reasoning plus alliance considerations create new failure modes.
| Model | Factual Accuracy | Spatial Accuracy |
| o1 | 94.29% | 100% |
| o1-mini | 56.33% | 84.93% |
| GPT-4o | 41.68% | 43.93% |
| Claude 3.5 Sonnet | - | 76.76% |
| DeepSeek-R1 | - | 97.38% |
@@
4.3.2 Factual Retrieval
Factual Retrieval in Classical Planning.
Here, we investigate whether LLMs can reliably retrieve key facts from a planning trajectory. For each task, we provide the model with the initial state (), the goal state (), and trajectories that explicitly detail the actions along with their corresponding state transitions. We control the number of mutable properties; for example, in a task involving moving balls, the ball location is considered mutable, whereas ball color might also be a property—but if no action can alter this property, we do not count it as mutable. We then prompt the model to report the state at a specified step. Figure 3 illustrates how retrieval accuracy varies with trajectory length. Notably, o1 performs most consistently, confirming that it ”reads” multi-step expansions more accurately than either GPT-4o or o1-mini.
Diplomacy Factual Knowledge Checks.
In Diplomacy, we design and categorize several factual queries into one-hop (e.g., ”Which locations does Russia control?”) vs. multi-hop (e.g., ”Which territories can France attack, and how many units of support are needed at least?”) to further check models’ factual retrieval in a highly strategic environment. As summarized in Figure 4, nearly all LLMs do well on basic location or adjacency checks but degrade by a large margin on ”Attackable” and ”Attack Analysis,” which demand deeper, multi-hop inference. Again, o1 and o1-preview lead, but still exhibit significant drops compared to simpler tasks.
4.3.3 Error Analysis
Error Analysis in Classical Planning
@@ 4.3.3 Error Analysis
To evaluate the propensity of various models to generate illegal moves during gameplay, and its ability to follow game constraint, we conduct analysis on three competitive games. Experimental settings and detailed results are shown in Appendix B.1. The result analysis underscores a clear correlation between the game complexity and the prevalence of illegal moves. Simpler games present minimal challenges, but in complex games(Chess), models exhibit a higher tendency to generate illegal moves to break the game constraint. This suggests that more sophisticated strategic planning and rule comprehension are required to excel in complex games.
@@ 4.4.1 Complex or Expansive Action Spaces
We further examine the results of 10 domains in classical planning tasks by analyzing two key measures: the average number of legal actions available at each step and the total state–action space, defined as the complete set of possible state–action pairs given an initial state and action space.To assess their impact on model performance, we compute the correlation between accuracy and these two measures. Surprisingly, as shown in Figure 5, accuracy exhibits a stronger correlation with the total state–action space rather than the average number of legal actions. This implies that cognitively, the model ”carries” the burden of large future branching factors, even if only a few choices are valid at each step.
@@
Turning to our solver vs. LLM matchups (Table 1 for main results, Appendix 8 for all results), solvers always win or draw the game. Tic-tac-toe reveals that advanced LLMs (e.g., o1, GPT-4-turbo, Claude 3.5 Sonnet) can achieve draws some of the time, but typically still lose or draw to the perfect solver. In Connect Four and Chess, the gap widens: our solver and Stockfish-level engines maintain a win rate across all tested LLMs. We also conduct analysis on the top k action percentages in Appendix(Figure 7). The distribution shows that while LLMs sometimes pick optimal moves in Connect Four, their accuracy drops drastically in Chess, underscoring how deeper tactics and branching expansions are beyond current LLMs’ capacity. Claude 3.5 Haiku and o3-mini fare somewhat better, yet remain far from human-expert or solver-level play.
Diplomacy Basic Skills.
Diplomacy features an even larger branching factor. As a baseline ”no negotiation” test, we pit one LLM against six neutral powers. There are only 3 out of 9 models pass the basic skill test: Table 2 shows that only o1-preview, o1 and GPT-4o can take 18 supply centers in the first 20 years. Other models time out after 20 years, reflecting the significant difficulty of long-horizon planning in such a vast environment. Details of each model and human’s performance on this basic skill test is shown in Appendix B.2.1.
@@ 4.4.2 Scaling Number of Agents
Hanabi with More Agents.
We vary the number of players from to in Hanabi and track final scores. Table 1 indicates that o1 remains the top performer, but its average score still declines from 16.4 (2-player) to 14.2 (5-player). This dip suggests that coordinating beliefs and actions among more agents still strains strong LLMs’ capacity to maintain consistent strategies over multiple incomplete information channels. However, models such as o1-mini and DeepSeek-R1 display erratic performance. Their high variance and relatively low average scores suggest insufficient strategic reasoning capabilities for effective Hanabi play.
Diplomacy with More Agents.
Diplomacy also allows variable numbers of participating powers. Detailed results of more multi-agent settings are shown in Appendix B.2.2. As the agent count grows (beyond 2-3 test seats for LLMs), we observe decreasing order accuracy, fewer successful attacks, and minimal supply-center gains. Ultimately, LLMs lose traction in highly interactive scenarios, underscoring how partial observability and shifting alliances further intensify the multi-agent complexity.
@@ 4.5 Social Intelligence and Negotiation Performance
Overall, we find that most current LLMs under-perform in both cooperative and negotiation-driven tasks, revealing critical gaps in their ability to manage dynamic, multi-agent contexts. In some cases (e.g., o1), large-scale social interaction even appears to negatively affect chain-of-thought coherence, underscoring an interplay between deep planning and social complexity that has not been thoroughly investigated in simpler benchmarks.
@@ 4.5.2 Negotiation in Diplomacy
Negotiation phases in Diplomacy—which theoretically enable players to form or dissolve alliances—often produce counterintuitive effects on strong planners like o1. As shown in Table 2 and Table 10, o1’s all kinds of order numbers drop when negotiation is introduced, whereas other models sometimes stay the same or even improve slightly. And final game scores Supply Centers and Controlled Regions for o1 also dropped by a large margin when enabling the negotiation, compared to other models. This result suggests that intense social interaction can disrupt planning coherence in otherwise capable LLMs, pointing to a tension between extended chain-of-thought reasoning and the cognitive overhead of real-time alliance-building, deception, or persuasion.
@@
To understand each model’s negotiation style in greater depth, we prompted o1 to annotate messages across six-agent Diplomacy settings using six LLM-assisted metrics (Table 2). We highlight several patterns: (Details of each metric is shown in Appendix A.4.2.)
@@
Persuasiveness: o1 exhibits the highest acceptance rate (), whereas Claude 3.5 Haiku is less effective ().
@@
Peace vs. Conflict Messaging: DeepSeek-R1 and Claude 3.5 Haiku demonstrate a relatively low peace ratio ( and ), showing their tendency to provoke conflicts, whereas o1 and GPT-4o strongly prefers peaceful approaches ( and ), showing conservatism in potentially adversarial negotiations.
@@
Perspective-Taking and Conditional Plans: Although moderate across the board, GPT-4-turbo leads in referencing other agents’ intentions (), while DeepSeek-R1 uses conditional ”if-then” tactics most frequently ().
Despite these varied negotiation styles, most LLMs—o1 included—lack flexibility and adaptivity akin to skilled human negotiators. Rather than shifting strategies in response to betrayals or emergent alliances, models often cling to simplistic patterns of communication, or in o1’s case, become overwhelmed by large volumes of social exchange.
These observations reinforce that contemporary LLMs exhibit only nascent social intelligence in scenarios where partial cooperation and deception can dramatically alter long-horizon outcomes.
@@ 5 Conclusion
Overall, our work highlights pressing gaps in both deep, multi-hop reasoning and robust social interaction, pointing toward the need for novel architectural innovations, integrated planning modules, and more advanced training methodologies. We hope SPIN-Bench will serve as a foundation and catalyst for continued progress in developing LLMs that are more strategically competent, socially adept, and ultimately closer to the vision of AGI.
@@
A primary limitation of this work lies in the use of pre-defined prompt templates. Additionally, our benchmark—though broad—cannot fully capture the complete spectrum of real-world strategic and social interactions. Consequently, the generalizability of our findings may be constrained. Future research could address these limitations by exploring more adaptive or automated prompt-engineering strategies and by extending the scope of SPIN-Bench to include an even wider array of tasks.
6 Impact Statements
Our work rigorously evaluated LLMs’ abilities in long-horizon decision-making, multi-agent coordination, and social reasoning. In doing so, SPIN-Bench provides a structured platform for both diagnosing current limitations and guiding the development of more advanced AI reasoning tools. Potential applications of this research span autonomous decision-making, game-theoretic analysis, and human–AI collaboration. However, these same capabilities raise important ethical concerns, particularly around fairness, manipulation, and unintended biases when highly capable systems engage in competitive or high-stakes settings. Ensuring that such systems remain transparent, accountable, and aligned with human values is a crucial aspect of responsible AI development. By identifying specific gaps in strategic and social competencies, this research not only advances the technical frontier but also highlights the need for frameworks that mitigate misuse and foster ethically grounded innovation.
Appendix A: Experiment Settings for Each Game
Appendix B: Supplementary Experiments
Appendix C: Prompt Design
Appendix A Experiment Settings
A.1 Classification and description of PDDL Domains

### v3

@@ 4.1 Outline and Setup
We evaluate a suite of popular LLMs that encompasses both closed-source commercial and open-source systems. The closed-source models include GPT-4o (Hurst et al., 2024), GPT-4o mini (OpenAI, 2024a), o1 (Jaech et al., 2024), o1-preview (OpenAI, 2024c), o1-mini (OpenAI, 2024b), o3-mini (OpenAI, 2025), GPT-4-turbo, GPT-3.5-turbo, Claude 3.5 Sonnet (Anthropic, 2024a), and Claude 3.5 Haiku (Anthropic, 2024b). Our open-source models include DeepSeek-R1 (Guo et al., 2025), Llama3-70b, Llama3.1-70b, Llama3.2-3b, Llama3.3-70b (AI, 2024a, b, c, d), Qwen2.5-72b (Team, 2024), and Mistral-7b (Jiang et al., 2023). Depending on the task, we select representative subsets to manage computational costs. Since Diplomacy exhibits the largest branching factor in our benchmark, we first conducted a basic skill evaluation to identify which models should advance to the full Diplomacy task. Detailed results for the basic skill evaluation are provided in Appendix C.2.1.
@@
We adopt multiple metrics suitable for each domain. In classical planning tasks, we use accuracy, N-Step Look Ahead, and plan validation through VAL (Planning, 2019). Competitive board games are tested via LLM vs. solver and LLM vs. LLM matches, Top-3 action percentages and we maintain internal Elo scores for leaderboards. In cooperative and strategic settings, we measure final scores, success rates for different action types, and advanced LLM-assisted negotiation metrics when negotiations are enabled. We excluded Diplomacy from the benchmark average score in Table 1 as few models play it well. Details of experimental settings, including prompt templates and domain rules, appear in Appendix B.
@@ 4.2 Overall Model Performance Across Tasks
As summarized in Tables 1, 2, the strongest closed-source models (e.g., o1, o1-mini, o3-mini) generally lead in classical planning, board-game play, and cooperative scenarios. Among open-source models, DeepSeek-R1 and higher-parameter Llama3 variants tend to outperform smaller-scale systems, although their margins behind the top-tier commercial LLMs remain significant.
@@ 4.3.1 Spatial Reasoning
To investigate whether LLMs’ planning deficits stem from weaker spatial understanding, we designed tasks requiring each model to track positions across sequences of relative movements. In these tasks, the LLM is given an initial spatial coordinate (e.g., (2,2)) along with a series of movements(up, down, left, right), and is tasked with computing the final coordinate. We evaluated all models on 90 scenes, each with a varying number of steps, to assess their spatial reasoning capabilities. Figure 2 plots the accuracy of each model against the length of the movement trajectory. Notably, o1-mini and GPT-4o exhibit declining performance as the number of steps increases, whereas o1 sustains perfect accuracy (100%) up to 29 steps.
4.3.2 Factual Retrieval
Factual Retrieval in Classical Planning.
We investigate whether LLMs can reliably retrieve key facts from a planning trajectory. For each task, we provide an initial state, the goal state, and trajectories that explicitly detail the actions along with their corresponding state transitions. We control the number of mutable properties; for example, in a task involving moving balls, the ball location is considered mutable, whereas ball color might also be a property—but if no action can alter this property, we do not count it as mutable. We then prompt the model to report the state at a specified step. Figure 3 illustrates how retrieval accuracy varies with trajectory length. Notably, o1 performs most consistently, confirming that it ”reads” multi-step expansions more accurately than either GPT-4o or o1-mini.
Diplomacy Factual Knowledge Checks.
In Diplomacy, we design and categorize several factual queries into one-hop (e.g., ”Which locations does Russia control?”) vs. multi-hop (e.g., ”Which territories can France attack, and how many units of support are needed at least?”) to further check models’ factual retrieval in a highly strategic environment. As summarized in Figure 4, nearly all LLMs do well on basic location or adjacency checks but degrade by a large margin on ”Attackable” and ”Attack Analysis,” which demand deeper, multi-hop inference.
4.3.3 Error Analysis
Error Analysis in Classical Planning
@@ 4.3.3 Error Analysis
| Claude 3.5 Haiku | 72.03 | 15.23 | 1.76 |
| DeepSeek R1 | 48.28 | 3.65 | 1.87 |
@@
To evaluate the propensity of various models to generate illegal moves during gameplay, and its ability to follow game constraint, we conduct analysis on three competitive games. Experimental settings and detailed results are shown in Appendix C.1. The result analysis underscores a clear correlation between the game complexity and the prevalence of illegal moves. Simpler games present minimal challenges, but in complex games(Chess), models exhibit a higher tendency to generate illegal moves to break the game constraint. This suggests that more sophisticated strategic planning and rule comprehension are required to excel in complex games.
Order Failure in Diplomacy.
We further examine action success rate in a more strategic setting, Diplomacy, where orders like Move and Attack require consistent tracking of boards and adjacency relations. Table 2 shows that o1 surpasses other models by on success rate of basic orders. However, all models’ performance drops markedly for complex moves requiring multi-step or multi-agent logic (e.g., Support-Self and Support-Others), suggesting that deeper reasoning plus alliance considerations create new failure modes.
@@ 4.4.1 Complex or Expansive Action Spaces
We further examine the results of 10 domains in classical planning tasks by analyzing two key measures: the average number of legal actions available at each step and the total state–action space, defined as the complete set of possible state–action pairs given an initial state and action space. To assess their impact on model performance, we compute the correlation between accuracy and these two measures. Surprisingly, as shown in Figure 5, accuracy exhibits a stronger negative correlation with the total state–action space rather than the average number of legal actions. This implies that cognitively, the model ”carries” the burden of large future branching factors, even if only a few choices are valid at each step.
@@
Turning to our solver vs. LLM matchups (Table 1 for main results, Appendix 8 for all results), solvers always win or draw the game. Tic-tac-toe reveals that advanced LLMs (e.g., o1, GPT-4-turbo, Claude 3.5 Sonnet) can achieve draws some of the time, but typically still lose or draw to the perfect solver. In Connect Four and Chess, the gap widens: all tested LLMs lose against the solver and Stockfish-level engines. We also conduct analysis on the top k action percentages in Appendix(Figure 7). The distribution shows that while LLMs sometimes pick optimal moves in Connect Four, their accuracy drops drastically in Chess, underscoring how deeper tactics and branching expansions are beyond current LLMs’ capacity.
| Model | Tic Tac Toe | Connect Four | Chess |
| Human | 1415 | 1692 | - |
| o1-preview | 1263 | 1377 | 1395 |
| o1-mini | 1205 | 1083 | 1247 |
| Claude 3.5 Sonnet | 1138 | 942 | 1196 |
| GPT-4o | 1025 | 980 | 1255 |
| Claude 3.5 Haiku | 983 | 907 | 1041 |
| Qwen2.5:72b | 944 | 877 | 1150 |
| Llama3.3 | 944 | 878 | 1146 |
| Llama3.1:70b | 936 | 829 | 1164 |
| GPT-4o-mini | 912 | 925 | 842 |
| GPT-4-turbo | 857 | 1007 | 1265 |
| Llama3:70b | 856 | 993 | 1155 |
| Mistral:7b | 855 | 800 | 243 |
| Llama3.2:3b | 836 | 821 | 421 |
| GPT-3.5-turbo | 829 | 890 | 480 |
Table 4 presents the internal Elo ratings of various LLMs. Among evaluated models, o1-preview and o1-mini lead in this category, demonstrating strongest strategic planning capabilities over the remaining models, further indicating the advantages of test time scaling in strategic games. Additionally, Claude 3.5 Sonnet, GPT-4o, and GPT-4-turbo maintain relatively high Elo ratings, indicating a superior ability to handle complex strategies compared to their counterparts. In contrast, models such as Mixtral, Llama3.2 perform significantly worse in Chess reflecting the increased demands that games with larger action and state spaces impose on a model’s planning and reasoning capabilities. This result underscores the current limitations of LLMs in strategic reasoning, highlighting the significant gap between AI and human-level planning capabilities.
@@ 4.4.2 Scaling Number of Agents
In Hanabi, we vary the number of players from to and track final scores. Table 1 indicates that o1 remains the top performer, but its average score still declines from 16.4 (2-player) to 14.2 (5-player). This dip suggests that coordinating beliefs and actions among more agents still strains strong LLMs’ capacity to maintain consistent strategies over multiple incomplete information channels. However, models such as o1-mini and DeepSeek-R1 display erratic performance. Their high variance and relatively low average scores suggest insufficient strategic reasoning capabilities for effective Hanabi play. In Diplomacy, we vary the agent number from 2 to 5. Detailed results of more multi-agent settings are shown in Appendix C.2.2. As the agent count grows (beyond 2-3 test seats for LLMs), we observe decreasing order accuracy, fewer successful attacks, and minimal supply-center gains. Ultimately, LLMs lose traction in highly interactive scenarios, underscoring how partial observability and shifting alliances further intensify the multi-agent complexity.
@@ 4.5 Social Intelligence and Negotiation Performance
Overall, we find that most current LLMs under-perform in both cooperative and negotiation-driven tasks, revealing critical gaps in their ability to manage dynamic, multi-agent contexts. In some cases (e.g., o1 and o1-preview), large-scale social interaction even appears to negatively affect chain-of-thought coherence, underscoring an interplay between deep planning and social complexity that has not been thoroughly investigated in simpler benchmarks.
@@ 4.5.2 Negotiation in Diplomacy
Negotiation phases in Diplomacy—which theoretically enable players to form or dissolve alliances—often produce counterintuitive effects on strong planners like o1, o1-preview. As shown in Table 2 and Table 10, o1’s all kinds of order numbers drop when negotiation is introduced, whereas other models sometimes stay the same or even improve slightly. And final game scores Supply Centers and Controlled Regions for o1, o1-preview also dropped by a large margin when enabling the negotiation, compared to other models. This result suggests that intense social interaction can disrupt planning coherence in otherwise capable LLMs, pointing to a tension between extended chain-of-thought reasoning and the cognitive overhead of real-time alliance-building, deception, or persuasion. In DeepSeek-R1, enabling negotiation features did not result in a significant performance decline. We attribute this to the model’s inherent tendency to propose one-sided solutions and provoke conflict during interactions (see the next section).
@@
We probe each model’s negotiation style in greater depth by prompting Claude 3.7 Sonnet to annotate messages across six-agent Diplomacy settings using six LLM-assisted metrics (see Appendix B.4.2 for details of the metrics; see Appendix C.2.3 for our human validation experiment of LLM annotation.). The results (see Table 2) enable the following key observations:
@@
Persuasiveness: o1 exhibits the highest acceptance rate (), whereas GPT-4o is less effective ().
@@
Peace vs. Conflict Messaging: DeepSeek-R1 and Claude 3.5 Haiku demonstrate a relatively low peace ratio ( and ), showing their tendency to provoke conflicts, whereas o1 strongly prefers peaceful approaches ().
@@
Perspective-Taking and Conditional Plans: Although moderate across the board, Claude 3.5 Haiku leads in referencing other agents’ intentions (), while DeepSeek-R1 uses conditional tactics most frequently ().
Despite these varied negotiation styles, most LLMs—o1 included—lack flexibility and adaptivity akin to skilled human negotiators. Rather than shifting strategies in response to betrayals or emergent alliances, models often cling to simplistic patterns of communication, or in o1’s case, become overwhelmed by large volumes of social exchange. These observations reinforce that contemporary LLMs exhibit only nascent social intelligence in scenarios where partial cooperation and deception can dramatically alter long-horizon outcomes.
@@ 5 Conclusion
Overall, our work highlights pressing gaps in both deep, multi-hop reasoning and robust social interaction, pointing toward the need for novel architectural innovations, integrated planning modules, and more advanced training methodologies. We hope SPIN-Bench will serve as a foundation and catalyst for continued progress in developing LLMs that are more strategically competent and socially adept.
@@
A primary limitation of this work lies in the use of pre-defined prompt templates. For the most complex environment Diplomacy, we conducted detailed ablation study on how prompt designs affect LLMs’ performance in Appendix D.3. But more detailed research is needed in future works. Additionally, our benchmark—though broad—cannot fully capture the complete spectrum of real-world strategic and social interactions. Consequently, the generalizability of our findings may be constrained. Future research could address these limitations by exploring more adaptive or automated prompt-engineering strategies and by extending the scope of SPIN-Bench to include an even wider array of tasks.
Appendix A: Reproducibility statement
Appendix B: Experiment settings for each game
Appendix C: Supplementary experiments
Appendix D: Prompt Design and Interaction Flow
Appendix A Reproducibility statement
We include a comprehensive setup and model specification in our code—scheduled for future release—to enable precise reproduction of our experimental results. But we also provide simple programs to run Tic Tac Toe, Connect Four, Chess, Hanabi and PDDL in our supplementary materials. Our package features scripts for running game engines, game solvers, the agent framework, and the entire benchmark arena. For commercial LLMs (e.g., Claude models), users can supply their own API keys. Additionally, to integrate new models into SPIN-Bench, we provide a generated PDDL dataset, solver scripts for three competitive games, multi-agent Diplomacy experiment settings, and all necessary interaction scripts.
Compute.
We conducted all evaluations on an Ubuntu 22.04 Linux machine equipped with 64 cores, 1024 GB of RAM, and 8 NVIDIA RTX 6000 GPUs (each with 50GB VRAM). For LLMs accessible only via endpoints—such as OpenAI and Claude—we utilized their commercial endpoints, thereby eliminating the need for local GPU resources. Additionally, we performed a detailed analysis of token usage across the various benchmark tasks. All results will be released alongside our code.
Appendix B Experiment Settings
B.1 Classification and description of PDDL Domains